# Enabling Vibration-Based Gesture Recognition on Everyday Furniture via Energy-Efficient FPGA Implementation of 1D Convolutional Networks

### v1

@@ I Introduction
The growing demand for smart home technologies has accelerated the development of interactive household interfaces, including vision-based, voice-controlled, and touch-enabled systems [1]. These systems enhance convenience by enabling intuitive control of lighting, climate, and appliances. However, conventional interaction methods rely on cameras, microphones, or capacitive sensors, each with inherent limitations. Camera-based approaches provide precise gesture and motion recognition but are sensitive to occlusion and lighting conditions, while also raising privacy concerns [2, 3]. Microphone-based systems enable hands-free use but are susceptible to ambient noise and pose privacy risks due to continuous audio recording [4, 5]. Capacitive sensors, widely used in touch panels, do not work reliably on thick or metallic surfaces, limiting their integration into furniture [6, 7].
To address these issues, Yoshida et al.[8] proposed Smatable (see Figure 1), a vibration-based sensing system that transforms ordinary tables into interactive surfaces. By attaching vibration sensors to the underside of tables, Smatable enables swipe gesture recognition without cameras or microphones, thereby improving privacy and material compatibility. Their follow-up work [9] improved accuracy by tuning Short-Time Fourier Transform (STFT) parameters for feature extraction and optimizing a 2D Convolutional Neural Network (2D-CNN). However, the enhanced model requires 1.38 GB of memory and relies on a costly high-performance CPU with high energy consumption. Their approach incurs a latency of 15 ms for STFT preprocessing and 350 ms for inference on an Intel Core i7-12700H, resulting in a total delay of 365 ms, well above the sub-100 ms threshold required for fluid interaction [10].
@@
We propose two compact 1D convolutional architectures tailored for embedded FPGAs: a standard 1D Convolutional Network (1D-CNN) and a Depthwise-Separable 1D Convolutional Network (1D-SepCNN). Compared to the 2D-CNN baseline [9] with over 369 million parameters, our models reduce the parameter count to just 216 while achieving comparable accuracy on the same datasets.
@@
We validate our solution on two swipe-direction datasets involving multiple users and tables. Under the PS data split setting, the selected 6-bit 1D-CNN achieves 0.970 accuracy with a 9.22 ms latency. The 8-bit 1D-SepCNN further reduces latency to 6.83 ms (over 53 speedup over CPU-based inference[9]) while maintaining comparable accuracy (0.949). Both models consume less than 1.2 mJ per inference, demonstrating strong suitability for energy-constrained edge deployment.
@@ II-A Vibration-based Interaction
Vibration-based sensing has emerged as a promising modality for enabling unobtrusive, material-agnostic, and privacy-preserving smart interfaces [11]. Early approaches typically relied on handcrafted features and traditional Machine Learning (ML) classifiers. For instance, SurfaceVibe [12] estimated wave propagation delays to detect taps and swipes, but suffered from limited adaptability due to fixed signal processing heuristics. VibSense [13] extracted time- and frequency-domain features and used classifiers such as Support Vector Machines and Random Forests, while iWood [14] embedded piezoelectric sensors into furniture and applied tree-based models for recognition. Although effective, these systems rely on manual feature design, limiting scalability and generalization.
To overcome these limitations, more recent work has turned to Deep Learning. Yoshida et al. [8, 9] demonstrated the feasibility of vibration-based gesture recognition using large 2D-CNNs applied to STFT-based spectrograms. While their system achieves high accuracy in controlled settings, it depends on computationally intensive preprocessing and high-capacity models, resulting in substantial latency and energy overhead. These drawbacks pose barriers to practical deployment on embedded devices. Our work addresses these challenges by rethinking the entire sensing-to-inference pipeline from a deployment perspective, aiming to achieve accurate, real-time recognition under stringent hardware constraints.
@@ II-B Compact Neural Architectures for Embedded Inference
To support deployment on constrained devices, NNs must strike a balance between recognition accuracy and computational efficiency. One often overlooked factor is the dimensionality of the input data. Larger inputs increase memory usage and amplify the computational burden in early network layers [15]. In vibration-based sensing, high-resolution inputs such as spectrograms contribute significantly to system overhead. Inspired by trends in speech and audio processing [16], we consider direct processing of downsampled raw waveforms, which can preserve essential temporal patterns while reducing input size and enabling real-time inference.
Architectural efficiency is equally critical. While prior work [9] adopts 2D-CNNs, several studies suggest that 1D-CNNs are better suited for sequential signals like vibration data [17, 18], offering a favorable trade-off between expressiveness and resource efficiency. Further improvements can be achieved through depthwise-separable convolutions, which have proven effective in embedded contexts [19]. These insights motivate our choice of both standard and separable 1D-CNN architectures tailored to the deployment constraints.
@@ II-C Quantization and FPGA Deployment
Beyond compact architectures, embedded deployment requires system-level optimizations that account for hardware constraints. Quantization is a widely adopted technique to reduce model size and computation by converting floating-point operations into low-bit integer arithmetic (e.g., INT8) [20, 21]. Integer-only inference is particularly advantageous on FPGAs, where custom datapaths can exploit reduced bitwidths to improve energy efficiency and throughput [22].
Moreover, hardware acceleration using FPGAs has shown great promise for achieving real-time, energy-efficient inference. For example, Rashid et al. [23] proposed TinyM2Net-V2, which leverages sparsity and hybrid quantization to accelerate inference on an Artix-7 FPGA. Liao et al. [24] introduced BearingPGA-Net, which uses parallel processing to speed up CNN-based fault detection on a Kintex-7 FPGA. While these works highlight the benefits of FPGA acceleration, they typically focus on high-performance FPGAs with ample computational resources. In contrast, fewer studies have investigated deployment on low-power embedded FPGAs such as the AMD Spartan-7, where resource budgets are much more constrained. These efforts underscore the need for end-to-end co-design strategies that align model structure, quantization, and hardware capabilities. Building on these insights, our work targets low-power FPGAs and explores quantization-aware model architecture and optimization techniques to enable efficient real-time gesture recognition using vibration data.
@@ III-A Onboard-Friendly Input Representation
Our study is based on the data collected by Yoshida et al.[9], which captures vibration signals from four synchronized piezoelectric sensors embedded in a table-mounted surface. The gestures consist of four swipe directions (Up, Down, Left, Right), sampled at 44.1 kHz and stored as 16-bit PCM waveforms. Two subsets are available: DataByPerson, where three users (A, B, C) perform gestures on the same table, and DataByTable, where person A interacts with three different tables (A, B, C) of varying height. Each gesture spans 2 seconds per sensor, totaling 352,800 samples per gesture. The original Smatable system [9] used STFT-based spectrograms (shape: 409690) as input to a large 2D-CNN, offering strong accuracy but incurring high compute and memory overhead.
@@ III-B Efficient Model Design
Building on the lightweight waveform input, we further improve deployability through efficient NNs design for embedded FPGAs. We develop two compact 1D convolutional architectures: a 1D-CNN and a depthwise-separable variant, 1D-SepCNN. Both architectures operate on downsampled waveforms and are optimized to reduce memory footprint and arithmetic complexity under hardware constraints. Table I presents an example configuration of 1D-CNN with three convolutional blocks (). In contrast to the 2D-CNN from [9], which contains 369 million parameters and operates on spectrograms of shape , our model directly processes downsampled waveforms of shape and contains only 296 parameters, achieving a drastic reduction in model size without significant loss in accuracy.
@@
To further reduce computational cost, we implement a variant using depthwise-separable convolution (SepConv1D). Each standard convolution is replaced by a DepthConv1D (applying a separate filter per input channel) followed by a PointConv1D (kernel size 1) that fuses information across channels [19]. This design substantially reduces parameter count and Floating-Point Operations (FLOPs). For example, replacing the Conv1D in the first block lowers parameters from 52 to 36 and FLOPs from 84,576 to 49,336. When applied to all three blocks, the total parameter count decreases to 216 (27.0% reduction), and FLOPs to 383,684 (33.1% reduction). These savings make 1D-SepCNN especially suitable for compute-bound FPGAs, where minimizing arithmetic workload is critical to achieving real-time performance.
@@ III-C Model Quantization and RTL Implementation
Building on the efficient model architectures, we apply quantization and RTL generation to enable integer-only inference on embedded FPGAs, where all numerical components (including weights, biases, and activations) and arithmetic operations are implemented using integer arithmetic [20]. This eliminates any floating-point (FP32) dependencies, significantly reducing hardware complexity and energy consumption [21]. We apply signed asymmetric quantization to both weights and activations to better utilize the available dynamic range. In contrast, bias terms and BatchNorm offsets are quantized symmetrically around zero for simplicity and hardware efficiency, as their distributions are typically centered and less sensitive to offset alignment.
This quantization scheme is supported by the open-source ElasticAI.Creator library222https://github.com/es-ude/elastic-ai.creator/tree/add-linear-quantization [25], which provides end-to-end support for QAT and integer-only inference across a wide range of NN components. The library offers synthesizable VHDL templates for standard components such as Dense layer, ReLU activation function, batch normalization, and global average pooling [26, 27].
@@ III-D Constraint-Aware Deployment Search
With the above optimizations in place, final deployment quality still hinges on selecting model configurations that balance accuracy and energy efficiency. To address this, we extend an open-source end-to-end deployment pipeline [22] to support the proposed 1D-CNN and 1D-SepCNN architectures. The framework builds on ElasticAI.Creator library and integrates a hardware-aware Optuna-based search framework to form a unified deployment flow from QAT to deployable FPGA implementations. As illustrated in Figure 4, it automates training, quantization, simulation, and profiling to guide the search toward viable configurations under practical constraints.
Unlike prior work [22] which prunes infeasible candidates based solely on resource usage, our framework applies early termination logic based on deployment-oriented criteria. Rather than completing all training and synthesis steps for every trial, constraint checks are applied progressively to prune invalid configurations as early as possible. This not only reduces redundant computation but also helps the Optuna sampler focus on promising regions of the search space. The constraint logic is fully implemented via Optuna’s trial pruning mechanism, without modifying its sampling strategy. The following constraint checks are incorporated:
@@ IV Results and Evaluations
The section starts by outlining the dataset splitting methods used in all experiments. We then reproduce the full-precision (FP32) 2D-CNN baselines from [9] under identical settings to establish fair references. Next, our constraint-aware, hardware-aware optimization framework is applied to search for efficient and deployable configurations of the proposed 1D-CNN and 1D-SepCNN models. Finally, we evaluate how well the selected configurations generalize across datasets.
@@ IV-A Datasets Splitting Methods
We adopt both the DataByPerson and DataByTable datasets333https://zenodo.org/records/17275491 from [9]. Each dataset comprises nine recording sessions per target (i.e., person or table). Within each session, four swipe directions are performed. For each direction, 10 distinct recordings are collected, resulting in 40 multichannel waveform recordings per session. As mentioned in Section III-A, we apply sliding-window augmentation with fractional offsets during downsampling, which expands each session from 40 to 400 recordings. We also adopt the data splitting methods in [9]: (1) Per-Subject (PS): Each target subject is split independently, with six sessions used for training and the remaining three for testing. (2) Leave-One-Subject-Out (LOSO): The model is trained on two targets and evaluated on the third, simulating cross-person or cross-table generalization. (3) Add-One-Session (AOS): One session from the held-out target is included in the training set, reflecting a practical scenario where limited target-specific data is available.
@@ IV-B Experiment 1: FP32 2D-CNN Baseline Reproduction
In [9], Yoshida et al. revised their 2D-CNN architecture and increased the number of STFT segments from 1000 to 4096, reporting improved classification accuracy compared to their earlier work [8]. However, only partial results were provided, specifically, test accuracies of 0.67 (LOSO) and 0.90 (AOS) on the DataByPerson dataset, averaged over three users. To establish a more comprehensive performance reference, we reproduce the FP32 2D-CNN model under the exact experimental settings described in their work and evaluate it across all three data splitting methods (PS, LOSO, AOS) on both the DataByPerson and DataByTable datasets.
Dataset | |||||||
| ByPerson | ByTable | |||||||||
| PS | A | A | 1.000 | 1.000 | ||||||
| B | B | 1.000 | 0.958 | |||||||
| C | C | 0.983 | 0.983 | |||||||
| avg. | 0.994 | 0.980 | ||||||||
| LOSO | B,C | A | 0.649 | 0.683 | ||||||
| A,C | B | 0.699 | 0.241 | |||||||
| A,B | C | 0.666 | 0.574 | |||||||
| avg. | 0.671 | 0.499 | ||||||||
| AOS | B,C + 1A | A | 0.941 | 0.850 | ||||||
| A,C + 1B | B | 0.800 | 0.733 | |||||||
| A,B + 1C | C | 0.949 | 0.941 | |||||||
| avg. | 0.897 | 0.841 | ||||||||
@@
As shown in Table II, our reproduced results on DataByPerson achieve average test accuracies of 0.994 (PS), 0.671 (LOSO), and 0.897 (AOS), closely matching the reported values in [9]. Minor deviations in significant digits are observed but remain within acceptable variance. On DataByTable, which was not evaluated in [9], the model performs well under PS (0.980), but shows significantly lower accuracy under LOSO (0.499) and moderate performance under AOS (0.841), indicating limited cross-table generalizability. Compared to their earlier work [8], these reproduced results yield mixed outcomes: occasionally improving accuracy, but often failing to generalize across different tables. These findings highlight the limitations of large, STFT-based 2D-CNN models in table-shifted scenarios, reinforcing the need for lightweight, robust alternatives that generalize better and support efficient deployment on resource-constrained platforms.
Model | Configuration | Test Accuracy |
| b | bs | lr ( | FP32 | Quantized | |||||||||||||||||||||||
| PS | 1D-CNN | 3 | 6 | 32 | 5.082 | 0.998 | 0.996 (0.20%) | 13.35 | 50.00 | 7.50 | 9.22 | 129 | 1.189 | ||||||||||||||
| 1D-SepCNN | 3 | 8 | 48 | 5.550 | 0.952 | 0.952 (0.00%) | 19.74 | 66.67 | 11.25 | 6.83 | 163 | 1.113 | |||||||||||||||
| LOSO | 1D-CNN | 5 | 6 | 56 | 3.251 | 0.744 | 0.738 (0.81%) | 18.98 | 73.33 | 10.00 | 20.94 | 152 | 3.182 | ||||||||||||||
| 1D-SepCNN | 3 | 6 | 56 | 3.330 | 0.702 | 0.675 (3.85%) | 16.08 | 50.00 | 11.25 | 6.83 | 146 | 0.996 | |||||||||||||||
| AOS | 1D-CNN | 4 | 8 | 32 | 4.810 | 0.948 | 0.941 (0.74%) | 19.08 | 83.33 | 8.75 | 13.32 | 159 | 2.118 | ||||||||||||||
| 1D-SepCNN | 5 | 8 | 48 | 9.967 | 0.930 | 0.909 (2.26%) | 29.82 | 96.67 | 16.25 | 11.16 | 199 | 2.221 | |||||||||||||||
@@
| = the number of convolutional blocks, = quantization bitwidth, = batch size, = learning rate. | |||||||||||||||||||||||||||
@@
| ∗ Latency measured on actual hardware deviates by 1.95% from simulation results. | |||||||||||||||||||||||||||
@@
| ∗∗ Power estimated from Vivado synthesis reports was validated on the actual FPGA hardware at 28.0°C, showing a deviation within 5.6%. | |||||||||||||||||||||||||||

### v2

@@ I Introduction
The growing demand for smart home technologies has accelerated the development of interactive household interfaces, including vision-based, voice-controlled, and touch-enabled systems [shi2022progress]. These systems enhance convenience by enabling intuitive control of lighting, climate, and appliances. However, conventional interaction methods rely on cameras, microphones, or capacitive sensors, each with inherent limitations. Camera-based approaches provide precise gesture and motion recognition but are sensitive to occlusion and lighting conditions, while also raising privacy concerns [camera1, camera2]. Microphone-based systems enable hands-free use but are susceptible to ambient noise and pose privacy risks due to continuous audio recording [microphone1, microphone2]. Capacitive sensors, widely used in touch panels, do not work reliably on thick or metallic surfaces, limiting their integration into furniture [capacitance1, capacitance2].
To address these issues, Yoshida et al.[yoshida2023smatable1] proposed Smatable (see Figure 1), a vibration-based sensing system that transforms ordinary tables into interactive surfaces. By attaching vibration sensors to the underside of tables, Smatable enables swipe gesture recognition without cameras or microphones, thereby improving privacy and material compatibility. Their follow-up work [yoshida2023smatable2] improved accuracy by tuning Short-Time Fourier Transform (STFT) parameters for feature extraction and optimizing a 2D Convolutional Neural Network (2D-CNN). However, the enhanced model requires 1.38 GB of memory and relies on a costly high-performance CPU with high energy consumption. Their approach incurs a latency of 15 ms for STFT preprocessing and 350 ms for inference on an Intel Core i7-12700H, resulting in a total delay of 365 ms, well above the sub-100 ms threshold required for fluid interaction [ng2012designing].
@@
We propose two compact 1D convolutional architectures tailored for embedded FPGAs: a standard 1D Convolutional Network (1D-CNN) and a Depthwise-Separable 1D Convolutional Network (1D-SepCNN). Compared to the 2D-CNN baseline [yoshida2023smatable2] with over 369 million parameters, our models reduce the parameter count to just 216 while achieving comparable accuracy on the same datasets.
@@
We validate our solution on two swipe-direction datasets involving multiple users and tables. Under the PS data split setting, the selected 6-bit 1D-CNN achieves 0.970 accuracy with a 9.22 ms latency. The 8-bit 1D-SepCNN further reduces latency to 6.83 ms (over 53 speedup over CPU-based inference[yoshida2023smatable2]) while maintaining comparable accuracy (0.949). Both models consume less than 1.2 mJ per inference, demonstrating strong suitability for energy-constrained edge deployment.
@@ II-A Vibration-based Interaction
Vibration-based sensing has emerged as a promising modality for enabling unobtrusive, material-agnostic, and privacy-preserving smart interfaces [huang2020vibration]. Early approaches typically relied on handcrafted features and traditional Machine Learning (ML) classifiers. For instance, SurfaceVibe [pan2017surfacevibe] estimated wave propagation delays to detect taps and swipes, but suffered from limited adaptability due to fixed signal processing heuristics. VibSense [liu2017vibsense] extracted time- and frequency-domain features and used classifiers such as Support Vector Machines and Random Forests, while iWood [wu2022iwood] embedded piezoelectric sensors into furniture and applied tree-based models for recognition. Although effective, these systems rely on manual feature design, limiting scalability and generalization.
To overcome these limitations, more recent work has turned to Deep Learning. Yoshida et al. [yoshida2023smatable1, yoshida2023smatable2] demonstrated the feasibility of vibration-based gesture recognition using large 2D-CNNs applied to STFT-based spectrograms. While their system achieves high accuracy in controlled settings, it depends on computationally intensive preprocessing and high-capacity models, resulting in substantial latency and energy overhead. These drawbacks pose barriers to practical deployment on embedded devices. Our work addresses these challenges by rethinking the entire sensing-to-inference pipeline from a deployment perspective, aiming to achieve accurate, real-time recognition under stringent hardware constraints.
@@ II-B Compact Neural Architectures for Embedded Inference
To support deployment on constrained devices, NNs must strike a balance between recognition accuracy and computational efficiency. One often overlooked factor is the dimensionality of the input data. Larger inputs increase memory usage and amplify the computational burden in early network layers [menghani2023efficient]. In vibration-based sensing, high-resolution inputs such as spectrograms contribute significantly to system overhead. Inspired by trends in speech and audio processing [nossier2020comparative], we consider direct processing of downsampled raw waveforms, which can preserve essential temporal patterns while reducing input size and enabling real-time inference.
Architectural efficiency is equally critical. While prior work [yoshida2023smatable2] adopts 2D-CNNs, several studies suggest that 1D-CNNs are better suited for sequential signals like vibration data [kiranyaz20191, shahid2022performance], offering a favorable trade-off between expressiveness and resource efficiency. Further improvements can be achieved through depthwise-separable convolutions, which have proven effective in embedded contexts [mo2021fpga]. These insights motivate our choice of both standard and separable 1D-CNN architectures tailored to the deployment constraints.
@@ II-C Quantization and FPGA Deployment
Beyond compact architectures, embedded deployment requires system-level optimizations that account for hardware constraints. Quantization is a widely adopted technique to reduce model size and computation by converting floating-point operations into low-bit integer arithmetic (e.g., INT8) [Jacob2017, Krishnamoorthi2018]. Integer-only inference is particularly advantageous on FPGAs, where custom datapaths can exploit reduced bitwidths to improve energy efficiency and throughput [ling2025deployment].
Moreover, hardware acceleration using FPGAs has shown great promise for achieving real-time, energy-efficient inference. For example, Rashid et al. [Rashid2024] proposed TinyM2Net-V2, which leverages sparsity and hybrid quantization to accelerate inference on an Artix-7 FPGA. Liao et al. [Liao2024] introduced BearingPGA-Net, which uses parallel processing to speed up CNN-based fault detection on a Kintex-7 FPGA. While these works highlight the benefits of FPGA acceleration, they typically focus on high-performance FPGAs with ample computational resources. In contrast, fewer studies have investigated deployment on low-power embedded FPGAs such as the AMD Spartan-7, where resource budgets are much more constrained. These efforts underscore the need for end-to-end co-design strategies that align model structure, quantization, and hardware capabilities. Building on these insights, our work targets low-power FPGAs and explores quantization-aware model architecture and optimization techniques to enable efficient real-time gesture recognition using vibration data.
@@ III-A Onboard-Friendly Input Representation
Our study is based on the data collected by Yoshida et al.[yoshida2023smatable2], which captures vibration signals from four synchronized piezoelectric sensors embedded in a table-mounted surface. The gestures consist of four swipe directions (Up, Down, Left, Right), sampled at 44.1 kHz and stored as 16-bit PCM waveforms. Two subsets are available: DataByPerson, where three users (A, B, C) perform gestures on the same table, and DataByTable, where person A interacts with three different tables (A, B, C) of varying height. Each gesture spans 2 seconds per sensor, totaling 352,800 samples per gesture. The original Smatable system [yoshida2023smatable2] used STFT-based spectrograms (shape: 409690) as input to a large 2D-CNN, offering strong accuracy but incurring high compute and memory overhead.
@@ III-B Efficient Model Design
Building on the lightweight waveform input, we further improve deployability through efficient NNs design for embedded FPGAs. We develop two compact 1D convolutional architectures: a 1D-CNN and a depthwise-separable variant, 1D-SepCNN. Both architectures operate on downsampled waveforms and are optimized to reduce memory footprint and arithmetic complexity under hardware constraints. Table I presents an example configuration of 1D-CNN with three convolutional blocks (). In contrast to the 2D-CNN from [yoshida2023smatable2], which contains 369 million parameters and operates on spectrograms of shape , our model directly processes downsampled waveforms of shape and contains only 296 parameters, achieving a drastic reduction in model size without significant loss in accuracy.
@@
To further reduce computational cost, we implement a variant using depthwise-separable convolution (SepConv1D). Each standard convolution is replaced by a DepthConv1D (applying a separate filter per input channel) followed by a PointConv1D (kernel size 1) that fuses information across channels [mo2021fpga]. This design substantially reduces parameter count and Floating-Point Operations (FLOPs). For example, replacing the Conv1D in the first block lowers parameters from 52 to 36 and FLOPs from 84,576 to 49,336. When applied to all three blocks, the total parameter count decreases to 216 (27.0% reduction), and FLOPs to 383,684 (33.1% reduction). These savings make 1D-SepCNN especially suitable for compute-bound FPGAs, where minimizing arithmetic workload is critical to achieving real-time performance.
@@ III-C Model Quantization and RTL Implementation
Building on the efficient model architectures, we apply quantization and RTL generation to enable integer-only inference on embedded FPGAs, where all numerical components (including weights, biases, and activations) and arithmetic operations are implemented using integer arithmetic [Jacob2017]. This eliminates any floating-point (FP32) dependencies, significantly reducing hardware complexity and energy consumption [Krishnamoorthi2018]. We apply signed asymmetric quantization to both weights and activations to better utilize the available dynamic range. In contrast, bias terms and BatchNorm offsets are quantized symmetrically around zero for simplicity and hardware efficiency, as their distributions are typically centered and less sensitive to offset alignment.
This quantization scheme is supported by the open-source ElasticAI.Creator library222https://github.com/es-ude/elastic-ai.creator/tree/add-linear-quantization [qian2023elasticai], which provides end-to-end support for QAT and integer-only inference across a wide range of NN components. The library offers synthesizable VHDL templates for standard components such as Dense layer, ReLU activation function, batch normalization, and global average pooling [ling2024flowprecision, ling2024integer].
@@ III-D Constraint-Aware Deployment Search
With the above optimizations in place, final deployment quality still hinges on selecting model configurations that balance accuracy and energy efficiency. To address this, we extend an open-source end-to-end deployment pipeline [ling2025deployment] to support the proposed 1D-CNN and 1D-SepCNN architectures. The framework builds on ElasticAI.Creator library and integrates a hardware-aware Optuna-based search framework to form a unified deployment flow from QAT to deployable FPGA implementations. As illustrated in Figure 4, it automates training, quantization, simulation, and profiling to guide the search toward viable configurations under practical constraints.
Unlike prior work [ling2025deployment] which prunes infeasible candidates based solely on resource usage, our framework applies early termination logic based on deployment-oriented criteria. Rather than completing all training and synthesis steps for every trial, constraint checks are applied progressively to prune invalid configurations as early as possible. This not only reduces redundant computation but also helps the Optuna sampler focus on promising regions of the search space. The constraint logic is fully implemented via Optuna’s trial pruning mechanism, without modifying its sampling strategy. The following constraint checks are incorporated:
@@ IV Results and Evaluations
The section starts by outlining the dataset splitting methods used in all experiments. We then reproduce the full-precision (FP32) 2D-CNN baselines from [yoshida2023smatable2] under identical settings to establish fair references. Next, our constraint-aware, hardware-aware optimization framework is applied to search for efficient and deployable configurations of the proposed 1D-CNN and 1D-SepCNN models. Finally, we evaluate how well the selected configurations generalize across datasets.
@@ IV-A Datasets Splitting Methods
We adopt both the DataByPerson and DataByTable datasets333https://zenodo.org/records/17275491 from [yoshida2023smatable2]. Each dataset comprises nine recording sessions per target (i.e., person or table). Within each session, four swipe directions are performed. For each direction, 10 distinct recordings are collected, resulting in 40 multichannel waveform recordings per session. As mentioned in Section III-A, we apply sliding-window augmentation with fractional offsets during downsampling, which expands each session from 40 to 400 recordings. We also adopt the data splitting methods in [yoshida2023smatable2]: (1) Per-Subject (PS): Each target subject is split independently, with six sessions used for training and the remaining three for testing. (2) Leave-One-Subject-Out (LOSO): The model is trained on two targets and evaluated on the third, simulating cross-person or cross-table generalization. (3) Add-One-Session (AOS): One session from the held-out target is included in the training set, reflecting a practical scenario where limited target-specific data is available.
@@ IV-B Experiment 1: FP32 2D-CNN Baseline Reproduction
In [yoshida2023smatable2], Yoshida et al. revised their 2D-CNN architecture and increased the number of STFT segments from 1000 to 4096, reporting improved classification accuracy compared to their earlier work [yoshida2023smatable1]. However, only partial results were provided, specifically, test accuracies of 0.67 (LOSO) and 0.90 (AOS) on the DataByPerson dataset, averaged over three users. To establish a more comprehensive performance reference, we reproduce the FP32 2D-CNN model under the exact experimental settings described in their work and evaluate it across all three data splitting methods (PS, LOSO, AOS) on both the DataByPerson and DataByTable datasets.
| Splitting Method | Training Data | Testing Data | Dataset | |
| ByPerson | ByTable | |||
| PS | A | A | 1.000 | 1.000 |
| B | B | 1.000 | 0.958 | |
| C | C | 0.983 | 0.983 | |
| avg. | 0.994 | 0.980 | ||
| LOSO | B,C | A | 0.649 | 0.683 |
| A,C | B | 0.699 | 0.241 | |
| A,B | C | 0.666 | 0.574 | |
| avg. | 0.671 | 0.499 | ||
| AOS | B,C + 1A | A | 0.941 | 0.850 |
| A,C + 1B | B | 0.800 | 0.733 | |
| A,B + 1C | C | 0.949 | 0.941 | |
| avg. | 0.897 | 0.841 | ||
@@
As shown in Table II, our reproduced results on DataByPerson achieve average test accuracies of 0.994 (PS), 0.671 (LOSO), and 0.897 (AOS), closely matching the reported values in [yoshida2023smatable2]. Minor deviations in significant digits are observed but remain within acceptable variance. On DataByTable, which was not evaluated in [yoshida2023smatable2], the model performs well under PS (0.980), but shows significantly lower accuracy under LOSO (0.499) and moderate performance under AOS (0.841), indicating limited cross-table generalizability. Compared to their earlier work [yoshida2023smatable1], these reproduced results yield mixed outcomes: occasionally improving accuracy, but often failing to generalize across different tables. These findings highlight the limitations of large, STFT-based 2D-CNN models in table-shifted scenarios, reinforcing the need for lightweight, robust alternatives that generalize better and support efficient deployment on resource-constrained platforms.
| Splitting Method | Model | Configuration | Test Accuracy | LUTs (%) | BRAMs (%) | DSPs (%) | Latency∗ (ms) | Power∗∗ (mW) | Energy (mJ) | ||||
| b | bs | lr ( | FP32 | Quantized | |||||||||
| PS | 1D-CNN | 3 | 6 | 32 | 5.082 | 0.998 | 0.996 (0.20%) | 13.35 | 50.00 | 7.50 | 9.22 | 129 | 1.189 |
| 1D-SepCNN | 3 | 8 | 48 | 5.550 | 0.952 | 0.952 (0.00%) | 19.74 | 66.67 | 11.25 | 6.83 | 163 | 1.113 | |
| LOSO | 1D-CNN | 5 | 6 | 56 | 3.251 | 0.744 | 0.738 (0.81%) | 18.98 | 73.33 | 10.00 | 20.94 | 152 | 3.182 |
| 1D-SepCNN | 3 | 6 | 56 | 3.330 | 0.702 | 0.675 (3.85%) | 16.08 | 50.00 | 11.25 | 6.83 | 146 | 0.996 | |
| AOS | 1D-CNN | 4 | 8 | 32 | 4.810 | 0.948 | 0.941 (0.74%) | 19.08 | 83.33 | 8.75 | 13.32 | 159 | 2.118 |
| 1D-SepCNN | 5 | 8 | 48 | 9.967 | 0.930 | 0.909 (2.26%) | 29.82 | 96.67 | 16.25 | 11.16 | 199 | 2.221 | |
@@
| = the number of convolutional blocks, = quantization bitwidth, = batch size, = learning rate. | |||||||||||||
@@
| ∗ Latency measured on actual hardware deviates by 1.95% from simulation results. | |||||||||||||
@@
| ∗∗ Power estimated from Vivado synthesis reports was validated on the actual FPGA hardware at 28.0°C, showing a deviation within 5.6%. | |||||||||||||